OpenStudy (anonymous):

Peggy had three times as many quarters as nickels. She had $1.60 in all. How many nickels and how many quarters did she have? Which of the following equations could be used to solve the problem? n + 3n = 1.60 n + 3n = 160 5n + 75n = 1.60 5n + 75n = 160

OpenStudy (anonymous):

Let q and n be the number of quarters and nickels, respectively. Solve the following for q and n:\[\left\{\frac{q}{n}=3,5 n+25 q=160\right\} \]\[\{n=2,q=6\} \]